OpenStudy (bbb911):

This is Geometry: Inga is starting a new rectangular garden. She wants the length to be 3 times the width, and she has 56 meters of fencing to put around the garden. Use the system of equations listed below to find the length "L" and width "w" of the garden. 2L+2w=56 and L +3w

OpenStudy (anonymous):

L + 3W = ?????

OpenStudy (bbb911):

wait sorry my mistake i mean L=3w

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

so L=3W ------------(1) 2L + 2W = 56 ------------(2) substitute the L in (2) by 3W 2(3W) + 2W = 56 8W = 56 W = 56/8 = 7

OpenStudy (anonymous):

so width = 7 L = 3(7) = 21

OpenStudy (bbb911):

Thank you so much :D

OpenStudy (anonymous):

welcome...